OpenStudy (australopithecus):

Find an equation of the tangent plane to the given surface at the specified point (2,-1,-3) for z =3y^(2) - 2x^(2) + x

OpenStudy (eyust707):

First we need to take partial derivatives

OpenStudy (eyust707):

Have you learned these

OpenStudy (australopithecus):

yes I know how to take partial derivatives

OpenStudy (eyust707):

\[f_{x} = -4x +1\] \(f_{y} = 6y\)

OpenStudy (australopithecus):

yes that is what I have gotten

OpenStudy (eyust707):

sweet

OpenStudy (australopithecus):

is the formula z-z_0 = f_x(x,y)(x-a) + f_y(x,y)(y-a)?

OpenStudy (australopithecus):

I assume it is not

OpenStudy (eyust707):

No you are correct! \[f_{x}(x_{0},y_{0})(x−x_{0})+f_{y}(x_{0},y_{0})(y−y_{0})−(z−z_{0})=0\]

OpenStudy (eyust707):

\(x_{0} = 2\) \(y_{0} = -1\) \(z_{0} = 3\)

OpenStudy (eyust707):

For the functions just plug in \(x_{0}\) and \(y_{0}\) into the partials.

OpenStudy (eyust707):

I got: \[-7(x−2)-6(y+1)−(z−3)=0\]

OpenStudy (eyust707):

you could simplify i suppose..

OpenStudy (australopithecus):

oh I see thanks

OpenStudy (eyust707):

no problem... kind of a silly formula to have to memorize.. i suppose you could learn to derive it but i never did lol

OpenStudy (australopithecus):

well it is similar to f'(x)(x-a) + f(x) = y from calc 1

OpenStudy (australopithecus):

so its not that hard to remember